Need ASAP
Just Answers

Question 36
Based on the capital asset pricing model, which one of the following must increase the expected return on an individual security, all else constant?

Select one:
a. An increase in the risk level of that security as measured by the standard deviation
b. An increase in the risk-free rate given a security beta of 1.42
c. A decrease in the market rate of return given a security beta of 1.13
d. A decrease in the market rate of return given a security beta of .78
e. A decrease in the risk-free rate given a security beta of 1.06


Question 37
Not yet answered
Marked out of 1.00
Not flaggedFlag question
Question text
Which one of the following is the vertical intercept of the security market line?

Select one:
a. Market rate of return
b. Individual security rate of return
c. Market risk premium
d. Individual security beta multiplied by the market risk premium
e. Risk-free rate


Question 38
Not yet answered
Marked out of 1.00
Not flaggedFlag question
Question text
You are assigned the task of computing the expected return on a portfolio containing several individual stocks. Which one of the following statements is correct concerning this task?

Select one:
a. The expected rate of return on the portfolio must be positive.
b. The arithmetic average of the betas for each security held in the portfolio must equal 1.0.
c. The portfolio beta must be 1.0.
d. The summation of the return deviation from the portfolio expected return for each economic state must equal zero.
e. The standard deviation of the portfolio must equal 1.0.


Question 39
Not yet answered
Marked out of 1.00
Not flaggedFlag question
Question text
BJB, Inc. stock has an expected return of 15.15 percent. The risk-free rate is 3.8 percent and the market risk premium is 8.6 percent. What is the stock's beta?

Select one:
a. 1.19
b. 1.21
c. 1.32
d. 1.48
e. 1.62


Question 40
Not yet answered
Marked out of 1.00
Not flaggedFlag question
Question text
You own a portfolio that has $1,900 invested in Stock A and $2,700 invested in Stock B. If the expected returns on these stocks are 9 percent and 15 percent, respectively, what is the expected return on the portfolio?

Select one:
a. 10.57 percent
b. 11.14 percent
c. 11.96 percent
d. 12.52 percent
e. 13.07 percent
Answer #1

36.
Correct answer is Option E (A decrease in the risk-free rate given a security beta of 1.06).

A decrease in the risk-free rate given a security beta of 1.06 should increase the expected return on an individual security, all else constant.

37.
Correct answer is Option E (Risk free rate).

Risk free rate is the vertical intercept of the security market line.

38.
Correct Answer is Option D.
The summation of the return deviation from the portfolio expected return for each economic state must equal zero.

39.
Expected return = Risk free rate + (Beta * Market risk premium)
15.15% = 3.8% (Beta * 8.6%)
Beta = (15.15% - 3.8%) / 8.6%
= 11.35% / 8.6%
= 1.32
Stock's beta = 1.32

40.
Beginning portfolio balance = 1,900 + 2,700 = $4,600.
Ending portfolio balance = (1,900 * 1.09) + (2,700 * 1.15) = $5,176

Expected return on portfolio = ($5,176 / $4,600) - 1
= 1.1252 - 1
= 0.1252 or 12.52%
Expected return on portfolio = 12.52%